# Approximating the order 2 quantum Wasserstein distance using the moment-SOS hierarchy

### v1

@@ 1 Introduction
Over the past decade there has been intensive efforts to generalize this theory to quantum settings where the density matrices replace the probability measures. A natural strategy for defining quantum analogues of optimal transport is the notion of quantum couplings, i.e., joint quantum states with fixed marginals. In analogy to the classical Monge–Kantorovich framework, this leads to defining transport costs via a cost operator $C$, where the total cost of transporting a density matrix $\rho$ to $\nu$ is given by
@@
defined over the set $Q(\rho,\nu)$ of possible joint quantum states, also called quantum couplings. However, this direct generalisation does not satisfy the desired Wasserstein distance properties. In [ZYYY22], the authors proved that there is no single cost operator that produces a sensible metric such as half the trace norm between states. Within this coupling-based paradigm, several alternative constructions have been proposed. The authors from [DPT21] construct couplings as density operators on a bigger Hilbert space with marginals $\rho^{T}$ and $\nu$. This formulation admits a canonical purification and naturally links couplings to quantum channels. The cost matrix is built from a set of self-adjoint operators (quadratures), yielding a Wasserstein-type distance $W^{S}_{2}(\rho,\sigma)$ through second-order moments. While $W^{S}_{2}$ is not a metric in the strict sense, it satisfies modified triangle inequalities and approximates classical Wasserstein distances in semiclassical regimes. The authors in [FECidZ22] use as their cost operator the projector onto the antisymmetric subspace of the tensor-product Hilbert space formed by two identical copies of the system. This construction gives rise to a quantity $W_{2}^{\mathrm{asym}}$ that captures the antisymmetric content of joint couplings. To recover data processing and monotonicity, a stabilized version $W_{2}^{\mathrm{asym},s}$ was introduced. For an overview of the various proposals for quantum Wasserstein distances and their properties, we refer the reader to the recent survey [Bea25].
@@
In this article we focus on the numerical implementation of $p=2,d=2$ case where the cost is quadratic in the pure states, and the quantum Wasserstein distance problem leads to an optimization problem over separable states– joint density oeprators with fixed marginals which can be expressed as a mixture of product states. This structure establishes the connection to the Doherty–Parrilo–Spedalieri (DPS) hierarchy [DPS02]. To test separability of quantum states, the DPS hierarchy provides a sequence of convex relaxations that impose symmetric extensions and positivity constraints that becomes increasingly tight at higher levels. Each relaxation bound can be computed by solving a semidefinite program [VB96], i.e., by minimizing a linear function under linear matrix inequality constraints.
In [GLS22], the authors provide a moment perspective for this hierarchy and show that state separability can be cast as an instance of the Generalized Moment Problem (GMP), which boils down to optimizing over the cone of positive Borel measures supported on the product of two unit spheres. Deeply inspired by this connection, we show that the order 2 quantum Wasserstein distance problem admits a similar GMP formulation, where one minimizes the squared Frobenius norm of the difference of rank-one projectors under marginal constraints. This allows us to use the moment-sums of squares (abbreviated as moment-SOS) hierarchy [Las01, Las09], initially designed to solve polynomial optimization problems. Each level of the hierarchy yields a lower bound on the true transport cost and the convergence is guaranteed. The resulting semidefinite relaxations are computationally tractable for moderate state dimensions and hierarchy levels.
@@ 2.1 Preliminaries on polynomial optimization
Let $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$ and $\overline{\mathbf{x}}=(\overline{x}_{1},\overline{x}_{2},\ldots,\overline{x}_{n})$ denote vectors of complex variables and their conjugates, respectively. Let $[\mathbf{x},\overline{\mathbf{x}}]$ denote the vector of monomials of the form $\mathbf{x}^{\alpha}\overline{\mathbf{x}}^{\beta}=\prod_{i}x_{i}^{\alpha_{i}}\prod_{j}\overline{x}_{j}^{\beta_{j}}$, where $\alpha,\beta\in\mathbb{N}^{n}$ are multi-indices. Let $[\mathbf{x},\overline{\mathbf{x}}]_{t}$ be the subvector consisting of monomials with total degree at most $t$, i.e., $|\alpha|+|\beta|\leq t$. We define the space of polynomials with complex coefficients and degree at most $t$ as,
@@
Given a polynomial $p=\sum_{\alpha,\beta}p_{\alpha,\beta}\mathbf{x}^{\alpha}\overline{\mathbf{x}}^{\beta}$, it conjugate is $\overline{p}=\sum_{\alpha,\beta}\overline{p}_{\alpha,\beta}\overline{\mathbf{x}}^{\alpha}\mathbf{x}^{\beta}$. A polynomial $p\in\mathbb{C}[\mathbf{x},\overline{\mathbf{x}}]$ is called Hermitian if $p=\overline{p}$. Let us denote by $\Sigma[\mathbf{x},\overline{\mathbf{x}}]$ the cone of Hermitian sums of squares of complex polynomials, and by $\Sigma[\mathbf{x},\overline{\mathbf{x}}]_{2t}$ its subset of elements with degree at most $2t$, for all $t\in\mathbb{N}$.
@@
[…]
@@ 2.2 Order 2 quantum Wasserstein distance
In the pure to pure case from Example 4, the transport plan is unique and yields an exact evaluation of the Wasserstein distance $W_{2}(\rho,\nu)=1$. This is in contrast to the mixed to mixed case from Example 5, where the provided feasible plan is not necessarily optimal and gives only an upper bound $1\geq W_{2}(\rho,\nu)$.
@@ 3.1 An instance of the Generalized Moment Problem
In a very similar way as in [GLS22, Section 5.2], we first show that the order 2 quantum Wasserstein distance can be formulated as a GMP instance. Let $\mathbf{X}=\{(\mathbf{x},\mathbf{y})\in\mathbb{C}^{2n}:\|\mathbf{x}\|=\|\mathbf{y}\|=1\}$ be the complex bi-sphere, i.e., the Cartesian product of two complex unit spheres, and let $\mathcal{M}(\mathbf{X})$ denote the set of finite positive Borel measures supported on $\mathbf{X}$. Let us note $f(\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}})=\mathrm{Tr}\left[(\mathbf{x}\mathbf{x}^{*}-\mathbf{y}\mathbf{y}^{*})\,(\overline{\mathbf{x}\mathbf{x}^{*}-\mathbf{y}\mathbf{y}^{*}})\right]\in\mathbb{C}[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]$ and consider the following GMP instance:
@@
The real reformulation is obtained similarly to [GLS22, Appendix A], where the authors provide detailed explaination to tranform polynomials in $\mathbb{C}[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]$ into polynomials in $\mathbb{R}[\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}]$, after applying the change of variables $\mathbf{x}=\mathbf{a}+\mathbf{i}\,\mathbf{b}$, $\mathbf{y}=\mathbf{c}+\mathbf{i}\,\mathbf{d}$. For any $p\in\mathbb{C}[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]$, let us define $p_{\operatorname{Re}}(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}):=\operatorname{Re}(p(\mathbf{a}+\mathbf{i}\,\mathbf{b},\mathbf{a}-\mathbf{i}\,\mathbf{b},\mathbf{c}+\mathbf{i}\,\mathbf{d},\mathbf{c}-\mathbf{i}\,\mathbf{d}))$. For a given Hermitian $p$ one has $p(\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}})=p_{\operatorname{Re}}(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d})$, so $f(\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}})=f_{\operatorname{Re}}(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d})$. The real analog of $\mathbf{X}$ is the set $\mathbf{X}_{\operatorname{Re}}=\{(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d})\in\mathbb{R}^{4n}:1-\|\mathbf{a}\|^{2}-\|\mathbf{b}\|^{2}=1-\|\mathbf{c}\|^{2}-\|\mathbf{d}\|^{2}=0\}$.
@@
  • The GMP instance (4) has a non-empty feasible set because $Q(\rho,\nu)$ is non-empty: for any spectral decompositions $\rho=\sum_{\ell}\omega_{\ell}u_{\ell}u_{\ell}^{*}$ and $\nu=\sum_{j}\omega^{\prime}_{j}v_{j}v_{j}^{*}$, one can obtain a finite sequence $\{(\omega_{\ell}\omega^{\prime}_{j},u_{\ell},v_{j})\}_{\ell,j}\in Q(\rho,\nu)$.
@@
    $\int_{\mathbf{X}}\operatorname{Tr}(\mathbf{x}\mathbf{x}^{*})\text{d}\mu=\int_{\mathbf{X}}\|\mathbf{x}\|^{2}\text{d}\mu=\int_{\mathbf{X}}\text{d}\mu=\operatorname{Tr}(\rho)=1.$
Hence, the inf of the real GMP instance (4) is attained at a finite atomic measure, as a consequence of Tchakaloff’s theorem [Tch57]. Equivalently, the inf of the complex GMP instance (2) is also attained at a probability measure $\mu=\sum_{\ell}\omega_{\ell}\delta_{(u_{\ell},v_{\ell})}$. This proves that $\{(\omega_{\ell},u_{\ell},v_{\ell})\}_{\ell}\in Q(\rho,\nu)$, and thus $W_{2}^{2}(\rho,\nu)\leq P(\rho,\nu)$.
@@
is weak-$\star$ closed. So the hypotheses of [Bar02, Theorem 7.2, Chapter 4] are satisfied and thus we have no duality gap between (2) and (3).
Remark 9.
@@
We emphasize that the dual program (3) might not be attained since the feasible set of (3) is not compact. For any feasible solution $(\Lambda,\Gamma)$ of (3), the tuple $(\Lambda+\alpha I,\Gamma-\alpha I)$ is also feasible for any arbitrary large $\alpha\in\mathbb{R}$ with the same objective value. Indeed, since $\operatorname{Tr}(\rho)=\operatorname{Tr}(\nu)=1$, one has $\operatorname{Tr}(\rho(\Lambda+\alpha I)+(\nu\Gamma-\alpha I))=\operatorname{Tr}(\rho\Lambda+\nu\Gamma)+\alpha\operatorname{Tr}(\rho-\nu)=\operatorname{Tr}(\rho\Lambda+\nu\Gamma)$. In addition, $\mathbf{x}^{*}(\Lambda+\alpha I)\mathbf{x}+\mathbf{y}^{*}(\Gamma-\alpha I)\mathbf{y}=\mathbf{x}^{*}\Lambda\mathbf{x}+\mathbf{y}^{*}\Gamma\mathbf{y}+\alpha(\|x\|^{2}-\|y\|^{2})=\mathbf{x}^{*}\Lambda\mathbf{x}+\mathbf{y}^{*}\Gamma\mathbf{y}$, for all $(x,y)\in\mathbf{X}$. The above observation implies that the feasible set of (3) is not bounded.

### v2

@@ 1 Introduction
Over the past decade there has been intensive efforts to generalize this theory to quantum settings where the density matrices replace the probability measures. A natural strategy for defining quantum analogues of optimal transport is the notion of quantum couplings, i.e., joint quantum states with fixed marginals. We refer to [FGP23] for a basic recent literature review. In analogy to the classical Monge–Kantorovich framework, this leads to defining transport costs via a cost operator $C$, where the total cost of transporting a density matrix $\rho$ to $\nu$ is given by
@@
defined over the set $Q(\rho,\nu)$ of possible joint quantum states, also called quantum couplings. However, this direct generalisation does not satisfy the desired Wasserstein distance properties. In [ZYYY22], the authors proved that there is no single cost operator that produces a sensible metric such as half the trace norm between states. Within this coupling-based paradigm, several alternative constructions have been proposed. The authors from [DPT21] construct couplings as density operators on a bigger Hilbert space with marginals $\rho^{T}$ and $\nu$. This formulation admits a canonical purification and naturally links couplings to quantum channels. The cost matrix is built from a set of self-adjoint operators (quadratures), yielding a Wasserstein-type distance $W^{S}_{2}(\rho,\sigma)$ through second-order moments. While $W^{S}_{2}$ is not a metric in the strict sense, it satisfies modified triangle inequalities and approximates classical Wasserstein distances in semiclassical regimes. The authors in [FECidZ22] use as their cost operator the projector onto the antisymmetric subspace of the tensor-product Hilbert space formed by two identical copies of the system. This construction gives rise to a quantity $W_{2}^{\mathrm{asym}}$ that captures the antisymmetric content of joint couplings. To recover data processing and monotonicity, a stabilized version $W_{2}^{\mathrm{asym},s}$ was introduced. As in the classical case, one can employ convex regularization [CGMP24] or von Neumann entropy regularization [GM23] for unbalanced noncommutative optimal transport problems. For an overview of the various proposals for quantum Wasserstein distances and their properties, we refer the reader to the recent survey [Bea25].
@@
In this article we focus on the numerical implementation of $p=2,d=2$ case where the cost is quadratic in the pure states, and the quantum Wasserstein distance problem leads to an optimization problem over separable states– joint density operators with fixed marginals which can be expressed as a mixture of product states. This structure establishes the connection to the Doherty–Parrilo–Spedalieri (DPS) hierarchy [DPS02]. To test separability of quantum states, the DPS hierarchy provides a sequence of convex relaxations that impose symmetric extensions and positivity constraints that becomes increasingly tight at higher levels. Each relaxation bound can be computed by solving a semidefinite program [VB96], i.e., by minimizing a linear function under linear matrix inequality constraints.
In [GLS22], the authors provide a moment perspective for this hierarchy and show that state separability can be cast as an instance of the Generalized Moment Problem (GMP), which boils down to optimizing over the cone of positive Borel measures supported on the product of two unit spheres. We refer the interested reader to [Las08] for a comprehensive introduction on the GMP, as well as for the practical approach based on semidefinite relaxations. Deeply inspired by this connection, we show that the order 2 quantum Wasserstein distance problem admits a similar GMP formulation, where one minimizes the squared Frobenius norm of the difference of rank-one projectors under marginal constraints. This allows us to use the moment-sums of squares (abbreviated as moment-SOS) hierarchy [Las01, Las09], initially designed to solve polynomial optimization problems. Each level of the hierarchy yields a lower bound on the true transport cost and the convergence is guaranteed. The resulting semidefinite relaxations are computationally tractable for moderate state dimensions and hierarchy levels.
@@ 2.1 Preliminaries on polynomial optimization
Let $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$ and $\overline{\mathbf{x}}=(\overline{x}_{1},\overline{x}_{2},\ldots,\overline{x}_{n})$ denote vectors of complex variables and their conjugates, respectively. Let $\mathbf{x},\overline{\mathbf{x}}]$ denote the vector of monomials of the form $\mathbf{x}^{\alpha}\overline{\mathbf{x}}^{\beta}=\prod_{i}x_{i}^{\alpha_{i}}\prod_{j}\overline{x}_{j}^{\beta_{j}}$, where $\alpha,\beta\in\mathbb{N}^{n}$ are multi-indices. Let $\mathbf{x},\overline{\mathbf{x}}]_{t}$ be the subvector consisting of monomials with total degree at most $t$, i.e., $|\alpha|+|\beta|\leq t$. We define the space of polynomials with complex coefficients and degree at most $t$ as,
@@
Given a polynomial $p=\sum_{\alpha,\beta}p_{\alpha,\beta}\mathbf{x}^{\alpha}\overline{\mathbf{x}}^{\beta}$, it conjugate is $\overline{p}=\sum_{\alpha,\beta}\overline{p}_{\alpha,\beta}\overline{\mathbf{x}}^{\alpha}\mathbf{x}^{\beta}$. A polynomial $p\in\mathbb{C}[\mathbf{x},\overline{\mathbf{x}}]$ is called Hermitian if $p=\overline{p}$. A Hermitian square is any complex polynomial of the form $\overline{p}\,p$, and any finite sum of such elements is called a sum of Hermitian squares. Let us denote by $\Sigma[\mathbf{x},\overline{\mathbf{x}}]$ the cone of sums of Hermitian squares of complex polynomials, and by $\Sigma[\mathbf{x},\overline{\mathbf{x}}]_{2t}$ its subset of elements with degree at most $2t$, for all $t\in\mathbb{N}$.
@@
    $L$ has a representing measure $\mu$ supported on $\mathbf{X}$, i.e., we have $L(p)=\int pd\mu$ for all $p\in\mathbb{C}[\mathbf{x},\overline{\mathbf{x}}]$.
@@ 2.2 Order 2 quantum Wasserstein distance
When the source and target are pure states, there is only one feasible transport plan. In the pure to pure case from Example 4, the transport plan is unique and yields an exact evaluation of the Wasserstein distance $W_{2}(\rho,\nu)=1$. This is in contrast to the mixed to mixed case from Example 5, where the provided feasible plan is not necessarily optimal and gives only an upper bound $1\geq W_{2}(\rho,\nu)$.
@@ 3.1 An instance of the Generalized Moment Problem
In a very similar way as in [GLS22, Section 5.2], we first show that the order 2 quantum Wasserstein distance can be formulated as a GMP instance. Let $\mathbf{X}=\{(\mathbf{x},\mathbf{y})\in\mathbb{C}^{2n}:\|\mathbf{x}\|=\|\mathbf{y}\|=1\}$ be the complex bi-sphere, i.e., the Cartesian product of two complex unit spheres, and let $\mathcal{M}(\mathbf{X})$ denote the set of finite positive Borel measures supported on $\mathbf{X}$. Let us denote $f(\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}})=\mathrm{Tr}\left[(\mathbf{x}\mathbf{x}^{*}-\mathbf{y}\mathbf{y}^{*})\,(\overline{\mathbf{x}\mathbf{x}^{*}-\mathbf{y}\mathbf{y}^{*}})\right]\in\mathbb{C}[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]$ and consider the following GMP instance:
@@
The real reformulation is obtained similarly to [GLS22, Appendix A], where the authors provide detailed explanation to transform polynomials in $\mathbb{C}[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]$ into polynomials in $\mathbb{R}[\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}]$, after applying the change of variables $\mathbf{x}=\mathbf{a}+\mathbf{i}\,\mathbf{b}$, $\mathbf{y}=\mathbf{c}+\mathbf{i}\,\mathbf{d}$. For any $p\in\mathbb{C}[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]$, let us define $p_{\operatorname{Re}}(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}):=\operatorname{Re}(p(\mathbf{a}+\mathbf{i}\,\mathbf{b},\mathbf{a}-\mathbf{i}\,\mathbf{b},\mathbf{c}+\mathbf{i}\,\mathbf{d},\mathbf{c}-\mathbf{i}\,\mathbf{d}))$. For a given Hermitian $p$ one has $p(\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}})=p_{\operatorname{Re}}(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d})$, so $f(\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}})=f_{\operatorname{Re}}(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d})$. The real analog of $\mathbf{X}$ is the set $\mathbf{X}_{\operatorname{Re}}=\{(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d})\in\mathbb{R}^{4n}:1-\|\mathbf{a}\|^{2}-\|\mathbf{b}\|^{2}=1-\|\mathbf{c}\|^{2}-\|\mathbf{d}\|^{2}=0\}$.
@@
  • The GMP instance (4) has a non-empty feasible set because $Q(\rho,\nu)$ is non-empty: for any spectral decomposition $\rho=\sum_{\ell}\omega_{\ell}u_{\ell}u_{\ell}^{*}$ and $\nu=\sum_{j}\omega^{\prime}_{j}v_{j}v_{j}^{*}$, one can obtain a finite sequence $\{(\omega_{\ell}\omega^{\prime}_{j},u_{\ell},v_{j})\}_{\ell,j}\in Q(\rho,\nu)$.
@@
    $\color[rgb]{0,0,0}\int_{\mathbf{X}}\text{d}\mu=\int_{\mathbf{X}}\|\mathbf{x}\|^{2}\text{d}\mu=\int_{\mathbf{X}}\operatorname{Tr}(\mathbf{x}\mathbf{x}^{*})\text{d}\mu=\operatorname{Tr}(\rho)=1.$}
Hence, the inf of the real GMP instance (4) is attained at a finite atomic measure, as a consequence of Tchakaloff’s theorem [Tch57]. Equivalently, the inf of the complex GMP instance (2) is also attained at a probability measure $\mu=\sum_{\ell}\omega_{\ell}\delta_{(u_{\ell},v_{\ell})}$. This proves that $\{(\omega_{\ell},u_{\ell},v_{\ell})\}_{\ell}\in Q(\rho,\nu)$, and thus $W_{2}^{2}(\rho,\nu)\leq P(\rho,\nu)$. Consequently, any optimal solution to the GMP instance (2) yields an optimal transport plan for the order 2 quantum Wasserstein distance problem.
@@
is weak-$\star$ closed. So the hypotheses of [Bar02, Theorem 7.2, Chapter 4] are satisfied and thus we have no duality gap between (2) and (3). ∎
Eventually, we show that the dual value $D(\rho,\nu)$ is attained if both source and target are positive definite quantum states.
Theorem 9.
Let $\rho,\nu\in\mathcal{D}(\mathbb{C}^{n})$ and $f$ as in Theorem 8. If both $\rho$ and $\nu$ are positive definite then the dual value $D(\rho,\nu)$ is attained.
Proof.
The proof follows from a compactness argument. Given a Hermitian matrix $\mathbf{A}\in\mathbb{C}^{n\times n}$, let us denote by $\|\mathbf{A}\|=\sqrt{\operatorname{Tr}(\mathbf{A}^{2})}$ the Frobenius norm of $\mathbf{A}$, by $\lambda_{\min}(\mathbf{A})$ and $\lambda_{\max}(\mathbf{A})$ the minimal and maximal eigenvalues of $\mathbf{A}$, respectively. Let us introduce the cone
$K=\left\{\left(\int\mathbf{x}\mathbf{x}^{*}\,\text{d}\mu,\int\mathbf{y}\mathbf{y}^{*}\,\text{d}\mu\right):\mu\in\mathcal{M}(\mathbf{X})\right\},$
whose dual cone $K^{*}$ is given by
$K^{*}=\bigl\{(\mathbf{A},\mathbf{B})\in\mathbb{C}^{n\times n}\times\mathbb{C}^{n\times n}\;:\;\mathbf{A}=\mathbf{A}^{*},\mathbf{B}=\mathbf{B}^{*},\operatorname{Tr}(\rho\mathbf{A}+\nu\mathbf{B})\geq 0\;\;\forall(\rho,\nu)\in K\bigr\}.$
Equivalently, using the definition of $K$, one can write
$K^{*}=\{(\mathbf{A},\mathbf{B})\in\mathbb{C}^{n\times n}\times\mathbb{C}^{n\times n}:\mathbf{x}^{*}\mathbf{A}\mathbf{x}+\mathbf{y}^{*}\mathbf{B}\mathbf{y}\geq 0\quad\forall(\mathbf{x},\mathbf{y})\in\mathbf{X}\}.$
Let $\varepsilon:=\min\{\lambda_{\min}(\rho),\lambda_{\min}(\nu)\}>0$. Since both $\rho-\varepsilon I$ and $\nu-\varepsilon I$ are positive semidefinite, there exists an atomic measure $\mu\in\mathcal{M}(\mathbf{X})$ such that $\int\mathbf{x}\mathbf{x}^{*}\,\text{d}\mu=\rho-\varepsilon I$ and $\int\mathbf{y}\mathbf{y}^{*}\,\text{d}\mu=\nu-\varepsilon I$, i.e., $(\rho-\varepsilon I,\nu-\varepsilon I)\in K$. Let $(\mathbf{A},\mathbf{B})\in K^{*}$. Duality implies that $\operatorname{Tr}((\rho-\varepsilon I)\mathbf{A}+(\nu-\varepsilon I)\mathbf{B})\geq 0$, i.e., $\operatorname{Tr}(\rho\mathbf{A}+\nu\mathbf{B})\geq\varepsilon\operatorname{Tr}(\mathbf{A}+\mathbf{B})$. From the equivalent definition of $K^{*}$, notice that $(\mathbf{A}+\mathbf{B})$ must be in particular positive semidefinite, since $\mathbf{x}^{*}\mathbf{A}\mathbf{x}+\mathbf{y}^{*}\mathbf{B}\mathbf{y}\geq 0$, for all $\mathbf{x}$. Then one has
$\|\mathbf{A}+\mathbf{B}\|\leq\sqrt{n}\lambda_{\max}(\mathbf{A}+\mathbf{B})\leq\sqrt{n}\operatorname{Tr}(\mathbf{A}+\mathbf{B}),$
which yields $\operatorname{Tr}(\rho\mathbf{A}+\nu\mathbf{B})\geq\varepsilon/\sqrt{n}\|\mathbf{A}+\mathbf{B}\|$.
Let us now consider a feasible pair $(\Lambda,\Gamma)$ for (3) and assume that the associated cost function is greater than 0. Let $c:=\varepsilon/\sqrt{n}$. By feasibility, one has $(I-\Lambda,I-\Gamma)\in K^{*}$, since $\mathbf{x}^{*}(I-\Lambda)\mathbf{x}+\mathbf{y}^{*}(I-\Gamma)\mathbf{y}=2-\mathbf{x}^{*}\Lambda\mathbf{x}-\mathbf{y}^{*}\Gamma\mathbf{y}\geq 0\quad\forall(\mathbf{x},\mathbf{y})\in\mathbf{X}$. Therefore
$\operatorname{Tr}(\rho(I-\Lambda)+\nu(I-\Lambda))\geq c\|2I-\Lambda-\Gamma\|.$
In addition,
$\operatorname{Tr}(\rho(I-\Lambda)+\nu(I-\Gamma))=2-\operatorname{Tr}(\rho\Lambda+\nu\Gamma)\leq 2.$
Therefore, $\|2I-\Lambda-\Gamma\|\leq 2/c$, thus the Frobenius norm of $\Lambda+\Gamma$ is upper bounded by $M:=2/c+2\sqrt{n}$, and by Cauchy-Schwartz $|\operatorname{Tr}(\Lambda+\Gamma)|\leq\sqrt{n}M=:M^{\prime}>0$.
@@
Note that the tuple $(\Lambda+\alpha I,\Gamma-\alpha I)$ is also feasible for any $\alpha\in\mathbb{R}$ with the same objective value. Indeed, since $\operatorname{Tr}(\rho)=\operatorname{Tr}(\nu)=1$, one has
$\operatorname{Tr}((\rho\Lambda+\alpha I)+(\nu\Gamma-\alpha I))=\operatorname{Tr}(\rho\Lambda+\nu\Gamma)+\alpha\operatorname{Tr}(\rho-\nu)=\operatorname{Tr}(\rho\Lambda+\nu\Gamma).$
In addition one has for all $(\mathbf{x},\mathbf{y})\in\mathbf{X}$
$\mathbf{x}^{*}(\Lambda+\alpha I)\mathbf{x}+\mathbf{y}^{*}(\Gamma-\alpha I)\mathbf{y}=\mathbf{x}^{*}\Lambda\mathbf{x}+\mathbf{y}^{*}\Gamma\mathbf{y}+\alpha(\|\mathbf{x}\|^{2}-\|\mathbf{y}\|^{2})=\mathbf{x}^{*}\Lambda\mathbf{x}+\mathbf{y}^{*}\Gamma\mathbf{y}.$
Therefore, let us also assume without loss of generality that $\operatorname{Tr}(\Lambda)=0$, which shall allow one to bound all eigenvalues of $\Lambda$ and $\Gamma$.
One has $M^{\prime}\leq\operatorname{Tr}(\Gamma)=\operatorname{Tr}(\Lambda+\Gamma)\leq 0$. For the latter inequality we used that for all $\mathbf{x}$ one has $f(\mathbf{x},\overline{\mathbf{x}},\mathbf{x},\overline{\mathbf{x}})=0$, so $\mathbf{x}(\Lambda+\Gamma)\mathbf{x}^{*}\leq 0$. In addition one has
$\leq\left(\sum_{i=1}^{n}|x_{i}|^{2}\right)^{2}+\left(\sum_{i=1}^{n}|y_{i}|^{2}\right)^{2},$
which implies that $\mathbf{x}\Lambda\mathbf{x}^{*}+\mathbf{y}\Gamma\mathbf{y}^{*}\leq 2$, for all $(\mathbf{x},\mathbf{y})\in\mathbf{X}$, so in particular $\lambda_{\max}(\Lambda)+\lambda_{\max}(\Gamma)\leq 2$. Since $\operatorname{Tr}(\Lambda)=0$, one has $\lambda_{\min}(\Lambda)\geq-(n-1)\lambda_{\max}(\Lambda)$. Combining this with $\lambda_{\max}(\Gamma)\leq-\lambda_{\min}(\Lambda)$ and $\lambda_{\max}(\Lambda)+\lambda_{\max}(\Gamma)\leq 2$ yields $\lambda_{\max}(\Gamma)\leq 2(n-1)/n$. In addition,
$\lambda_{\min}(\Gamma)\geq-(n-1)\lambda_{\max}(\Gamma)+\operatorname{Tr}(\Gamma)\geq-2(n-1)^{2}/n-M^{\prime}.$
Eventually, one has $\lambda_{\max}(\Lambda)\leq-\lambda_{\min}(\Gamma)\leq 2(n-1)^{2}/n+M^{\prime}$ and $\lambda_{\min}(\Lambda)\geq-(n-1)\lambda_{\max}(\Lambda)\geq-2(n-1)^{3}/n-(n-1)M^{\prime}$. Consequently, all eigenvalues of such feasible matrix tuples are bounded in absolute value, which implies that all their entries’ absolute values are also bounded.
The feasible set (3) is the intersection of infinitely many polynomial superlevel sets. Therefore it is closed, which concludes the desired dual attainment proof. ∎